OpenStudy (anonymous):

I need help with this improper integral example please:

OpenStudy (anonymous):

\[\int\limits_{1}^{b}\frac{ 1 }{ x ^{2} }dx=\left[ -x ^{-1}\right]_{1}^{b}\]

OpenStudy (anonymous):

i dont get the step where I go from x^2 to -x^-1

hartnn (hartnn):

\(\huge \int x^n dx = \dfrac{x^{n+1}}{n+1}+c\) here 1/x^2 means x^-2 so, n= -2

OpenStudy (anonymous):

wow thanks a lot! really couldn't recall this, my brain froze. appreciate it :)

hartnn (hartnn):

welcome ^_^